a rectangle with 4cm and 10cm is revolved about its longer side,find the volume and surface area
Answer by Alan3354(69443)

a rectangle with 4cm and 10cm is revolved about its longer side,find the volume and surface area
---------------
Area = 4*10 = 40 cm^2
The centroid is 2 cm from the axis of rotation.
Theorem of Pappus --> Vol = 40*2*2pi = 160pi cubic cm.
------
The lateral area is that of a cylinder with r = 10 cm and h = 4 cm